The symmetry of (2,0,0) vibrational state of co2 is
1
Expert's answer
2019-05-06T10:14:05-0400

Carbon dioxide is a linear molecule.

The degrees of vibrational modes for linear moleculeі can be calculated using the formula:


"3 \times N – 5 (1)"

where N is the number of atoms in molecule

Carbon dioxide has

"3 \times 3 - 5 = 4 vibrations."


Answer:

4